Left Front Bent Gregorian Technical Observing with IRTC: UT 2010 January 27

Observer: DMiller (LBTO)
Telescope Operator: JHill
Instrument Support: DOfficer
Telescope Support: JLittle
SW support: TLeibold on call

Summary

Partly cloudy weather, but it isn't snowing!

Open from 01:00 until 05:48 for LFBG tests. Mostly trying to reproduce the non-convergence of Issue #2500.

Reopen from 07:30 until 10:25 for LBC-Blue tests to measure the LBCFPIA crosstalk terms between Z4-Z11-Z22.

Preparations

00:30 JMorris performed telescope walk-through. JLittle configured and balanced the telescope for LBFG.

Details

All Times in UT

Open

01:00 JLittle started telescope and opened chamber

Rotator Problem

01:20 JHill notes that the rotators are not available. Calls TSargent, then DAshby and finally DCox.

01:40 Error 1 or both rotators doesn't exist. (after rot -f command)

Dan give instruction to switch a soft-link on jet at /home/dsp/dps.conf from /home/dsp/dg.conf to /home/dsp/lucifer.conf

01:50 Rotator is now communicating. Note that Dan and Tom were able to fix this problem the next day and close Issue #2467.

01:55 IRTC at LFBG is authorized

02:00 Start IRAF and register IRTC@LFBG

02:01 pointto BS9126 TelMode=ACTIVE gs=0 Failed, o

Pointing Model Problem

02:01 IRTC at LFBG is authorized again. John notes that both authorizes start but not return

02:05 Chris does not answer phone. Torsten is called.

02:09 Stop and start IIF. Still does not work

02:14 Torsten finds that PCS is not returning to IIF after the de-authorize? command

02:15 Michele is called. She looked through the log and calls us back. PCS was having a "waiting for handle block" or something.

02:20 Michele calls and had us restart PCS. IRTC@LFBG authorized. Success.

02:25 pointto sent again. MCS is now "waiting for handle block". Killed MCS and restarted. We think these two hung subsystems may have been caused by stopping and restarting MCSPU earlier.

02:34 pointto BS9126 TelMode=ACQUIRE gs=0 Success. The acquisition image looks a bit odd. There is a dark band along the bottom and right side of image. Moonlight?

02:39 pointto M5_0269 TelMode=ACQUIRE gs=0 Success

02:41 set IA to -400 Star is seen just off the guider lower left

02:43 John sets IE = -25 CA -8

02:44 pointto BS9126 TelMode=ACTIVE gs=0 Success. Collimating out a bunch of coma. M1 is 5C too warm. Collimated to 400 nm rms to wfsc 21. Seeing is ~1.5

03:02 Michele calls back and says there is a kernel problem cause by the pointing model (from 10200102). Tried to re-authorize as LBC. PCS died. Successfully loaded an old pointing model with no problem. Michele will investigate later. John tried to load the new pointing model and this time it was successful!!!

03:08 Authorize LUCIFER@LFBG because collimation model is better then IRTC one.

JMH replaced the old IRTCF collimation model with the 20100105 LUCIFER collimation model the following day.

03:10 pointto M5_0269 TelMode=ACQUIRE gs=0 PAM- See star at the top of image. IE=-58 CA=-1 at AZ=97 EL=51

03:13 pointto BS9169 Start collimate at wfsc 22.

Reflective Memory Problem ?

03:21 "external centrioding routine did not return" Guiding and Active Optics stopped. Chris thinks this was the reflective memory/network problem, and not any problem with guiding.

Hunting for the Issue 2500 Non-convergence at low-elevation

03:24 resend pointto BS919 Success. Collimate to 36

03:31 pointto BS9101 at Elevation 27 to test for collimation problem (Issue #2500)

wfsc rot angle Looks
37-45 310 Maybe a bit on the lower left
46-48 290 Maybe a bit on the left
49-50 330 OK
51 270 OK
52 130 OK
53 40 OK
54 0 Maybe
55-56 -30 OK
57-64 -60 Maybe

03:50 offsetpointing with rot=-110 from original: couldn't move probe to 0.5, -0.5 (or so) Refraction was suggested by John for the cause of trying to move probe past 612.5 (or so...)
Differential refraction between the guide image and the on-axis IRTC image can make the probe wan't to move out past the on-axis position depending on the location in the sky and the PA. This was Issuetrak'd later as Issue #2597.

03:51 pointto BS9101 TelMode=ACTIVE gs=0 PA=-90 Failed, couldn't move probe

03:51 pointto BS9101 TelMode=ACTIVE gs=0 PA=-180 probe moved, but star not found (off the side of acq image)

03:54 pointto BS9101 TelMode=ACQUIRE gs=0 PA=-180 John set IE=-80 CA=-8

04:03 pointto BS9101 TelMode=ACTIVE gs=0 PA=-180 Success

04:16 Down to 17 elevation

04:17 pointto BS9105 TelMode=ACTIVE gs=0 PA=-300 Success

wfsc rot angle Looks
65-76 17 OK
77-81 -44 OK

04:32 pointto BS9105 TelMode=ACTIVE gs=6 PA=337 Success

x=0, y=124

wfsc rot angle Looks
82-86 -67 OK

04:37 pointto BS9105 TelMode=ACTIVE gs=6 PA=320 Success

x=37, y=118

wfsc rot angle Looks
87-89 -85 OK

04:37 pointto BS9105 TelMode=ACTIVE gs=6 PA=320 Success

x=-26, y=121

wfsc rot angle Looks
90-93 -55 OK

04:45 John move telescope to el=80

04:48 pointto BS9105 TelMode=ACTIVE gs=6 PA=320 Success

x=-26, y=121 Elevation 34

wfsc rot angle Looks
94-95 -56 OK
96-97 -86 OK
98-108 -27 OK
109-110 -57 OK

05:04 John moves telescope to horizon

05:05 pointto BS9105 TelMode=ACTIVE gs=6 PA=350

x=-26, y=121 Elevation 34

Failed: no guide star on acq image

05:08 pointto BS9105 TelMode=ACQUIRE gs=0 PA=350 Set IE=-100 CA=-15

05:12 pointto BS9105 TelMode=ACTIVE gs=6 PA=350

wfsc rot angle Looks
111-121 -58 OK
122 -88 OK
123 -28 CLOUDS

05:30 pointto M5_0354 TelMode=ACTIVE gs=0 PA=-90 (Low in South East)

Primary reached collimation limit. John put -1.0 mm Global Y on Primary and Secondary

05:32 pointto BS9141 TelMode=ACTIVE gs=0 PA=-90 No star

Close

05:48 close for clouds and radar echoes

Determining the mounting positions for IRTC and LUCIFER

05:56 pointto BS9123 TelMode=ACQUIRE gs=0

Rotator angle 7 AZ=180.75

This set the rotator such that North is up on Lucifer and the +Y axis is anti-aligned with the PCS_XY +Y axis. Also, the rotator should be at the center of it's rotational travel. The first two alignments seem to be correct; the mounting orientation of Lucifer and the AGw are correct. However, the rotator is not in its center of rotational travel, but off my 180 degrees. The opening in the cable wrap should be positioned straight up. See attached picture.

Reconfigure of Telescope to LBC

John took about 25 minutes of the cloudy time to retract M2 and M3 and deploy LBC

07:25 Ready for LBC observing

Open

07:30 Re-open for LBC Tests in partly cloudy sky.

See nightreport for LBC filenames of test images. These are also listed in PartnerObserving.20100127UT.

Close

10:25 Close for clouds

-- DougMiller - 27 Jan 2010
Topic revision: r5 - 16 Mar 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback